OpenStudy (anonymous):

Which of the following shows the correct order of the numbers below? (1/2)^2, 2/3 , 0.45 (1/2)^2 < 0.45 < 2/3 (1/2)^2 < 2/3 < 0.45 0.45 < (1/2)^2 < 2/3

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (here_to_help15):

(1/2)^2 = 1/4 = 0.25 2/3 = 0.666

OpenStudy (anonymous):

so B?

OpenStudy (here_to_help15):

Yes its B lol

OpenStudy (anonymous):

thnx (I don't mind direct answers:P)

OpenStudy (cwrw238):

write the 3 of them done as decimals then its easy;- 0.45, 0.25, 0.67

OpenStudy (here_to_help15):

Something tells me your telling me that it's D 0.o

OpenStudy (cwrw238):

the order is 0.25, 0.45 , 0.67 - now convert them back to their original form in the question

OpenStudy (cwrw238):

look at here-to-help's post which converted them to decimals